### v1

# Spectroscopic Characterization of Interstellar Object 3I/ATLAS: Water Ice in the Coma

Abstract
We present optical and near-infrared spectroscopy of the interstellar object 3I/ATLAS, obtained with Gemini-S/GMOS and NASA IRTF/SpeX on 2025 July 5 and July 14. The optical spectrum shows a red slope of 10% per 1000 Å between 0.5 and 0.8 m, closely resembling that of typical D-type asteroids. At longer wavelengths, the near-infrared spectrum flattens significantly to 3% per 1000 Å from 0.9 to 1.5 m, consistent with the spectral behavior of large water ice grains in the coma. Spectral modeling with an areal mixture of 70% Tagish Lake meteorite and 30% 10 m-sized water ice successfully reproduces both the overall continuum and the broad absorption feature near 2.0 m. The 1.5 m water ice band, however, is not detected—likely due to the limited signal-to-noise of the IRTF data and dilution by refractory materials. The 30% ice fraction should be interpreted as an approximate, order-of-magnitude estimate of the coma composition. The agreement between the GMOS and SpeX spectra, taken nine days apart, suggests short-term stability in the coma’s optical properties. Our observations reveal that 3I/ATLAS is an active interstellar comet containing abundant water ice, with a dust composition more similar to D-type asteroids than to ultrared trans-Neptunian objects.
UTF8gbsn
1 Introduction
The third interstellar object (ISO) was discovered on 2025 July 1 by the ATLAS Chile telescope (J. L. Tonry et al., 2018) during its robotic sky survey and was initially designated A11pl3Z. This was a particularly challenging detection because the object lay in Sagittarius, near the galactic plane, a region often avoided by most surveys due to stellar crowding. Following the discovery, numerous observatories quickly conducted follow-up observations, and archival searches uncovered pre-discovery images that extended the orbital arc back to 2025 May 22. Early measurements indicated an orbital eccentricity greater than 6, and reports of coma or activity prompted its official designation as 3I/ATLAS = C/2025 N1 (ATLAS) (MPEC 2025 N12).
The initial characterization of 3I/ATLAS (D. Z. Seligman et al., 2025) highlighted both similarities and differences relative to the first two known interstellar objects. Whereas 1I/ʻOumuamua and 2I/Borisov were both estimated to have sub-kilometer radii (A. Fitzsimmons et al., 2024), early photometric measurements for 3I suggested a nucleus radius of less than 10 km. This estimate was derived from scattered light measurements that included contributions from both the nucleus and surrounding coma dust, assuming a typical cometary albedo. Unlike 1I/ʻOumuamua, which showed no detectable dust coma during its brief observation window, both 2I/Borisov and 3I/ATLAS were discovered with prominent dust comae. Nevertheless, 1I must also have been outgassing, as inferred from the non-gravitational acceleration detected in its trajectory (M. Micheli et al., 2018). Determining the nucleus rotation has been challenging due to the dense galactic plane background and the presence of a dust coma. However, one study reports a low-amplitude light curve with m 0.08 mag and a period near 17 hours (R. de la Fuente Marcos et al., 2025).
Spectral reflectance measurements of 3I consistently indicate red dust, with reported spectral slopes of = 17–19% per 1000 Å between 0.4–0.8 m (D. Z. Seligman et al., 2025; C. Opitom et al., 2025; M. Belyakov et al., 2025), similar to D-type asteroids, and a flattening of the slope beyond 0.8 m. Gas emissions from CN were detected for 2I/Borisov about a week after discovery (A. Fitzsimmons et al., 2019) at production rate of Q(CN) = 4 1024 molec/s when it was at =2.7 au. In contrast, to date, no gas emissions have been detected from 3I with an upper limit of Q(CN)4.91023 molec/s (A. Alvarez-Candal et al., 2025). However, this limit was obtained at 4.4 au where it is still quite cold to expect CN. Only a handful of comets have ever had CN detected beyond 4 au and on the order of a dozen beyond 3 au (M. F. A’Hearn et al., 1995; A. L. Cochran et al., 2012).
In this paper, we present spectroscopic observations of 3I/ATLAS aimed at characterizing its dust coma.
2 Observations and Data Reduction
2.1 Gemini
The Gemini Multi-Object Spectrograph (GMOS) on the Gemini South telescope was used to obtain low resolution spectra of 3I/ATLAS on 2025 July 4 and 5 UT. Triggering a rapid target of opportunity program [GS-2025A-Q-113], we obtained 6 exposures using the R150 grating and a 1.0′′ slit, with central wavelength at 0.6 m. We used a spatial binning of 22 and spectral binning 44. We also also observed two G2V solar analog stars (HD 174466 & HD 167221) at airmasses similar to 3I, serving as both telluric correction standards and as solar analog references. The data were reduced using the DRAGONS (Data Reduction for Astronomy from Gemini Observatory North and South) pipeline (K. Labrie et al., 2019), the official data reduction package provided by the Gemini Observatory.
| Observation Date | Tel./Inst. | Object | Exposure | Airmass | aaHeliocentric distance | bbGeocentric distance | ccPhase angle |
| (UT) | (s) | (au) | (au) | (∘) | |||
| 2025-07-04 | Gemini/GMOS | 3I | 6200 | 1.26–1.35 | 4.404 | 3.408 | 2.96 |
| HD 174466 | 43 | 1.10–1.11 | |||||
| 2025-07-05 | Gemini/GMOS | 3I | 6200 | 1.37–1.50 | 4.371 | 3.379 | 2.65 |
| HD 167221 | 43 | 1.23–1.25 | |||||
| HD 156734 | 815 | 1.46–1.48 | |||||
| 2025-07-14 | IRTF/SpeX | 3I | 28120 | 1.27–1.36 | 4.072 | 3.143 | 6.60 |
| HD 167221 | 1612 | 1.25–1.26 |
2.2 IRTF
Observations were conducted on the evening of 2025 July 14 (UT) using Director’s Discretionary Time (DDT) with the SpeX spectrograph on the NASA Infrared Telescope Facility (IRTF). We conducted observations using the SpeX instrument in combination with the MIT Optical Rapid Imaging System (MORIS) camera (A. A. S. Gulbis et al., 2011). We employed the 0.7 m dichroic in SpeX, which transmits near-infrared light to the spectrograph while reflecting visible light (m) to MORIS. Instead of using the facility guiding system (GuideDog), we guided directly on the target using MORIS. This setup enabled stable guiding of the faint interstellar object in a crowded field, while simultaneously acquiring near-infrared spectra with SpeX. NIR data were acquired in prism mode covering the 0.8–2.5 m wavelength range, using a 0.8′′ 15′′ slit, which provided a spectral resolving power of R150 (J. T. Rayner et al., 2003). Two nearby G-type solar analog star observed along with 3I, served both as the telluric correction standard and as a solar analog reference. The spectral data were reduced using the SpeXtool data reduction pipeline (M. C. Cushing et al., 2004). A journal of the observational details and target geometry is provided in Table 1.
3 Observational Analysis
The GMOS spectrum of 3I/ATLAS is shown in Figure 1. It is largely featureless and exhibits a moderately red spectral slope. The comet was observed on two nights, 2025 July 4 and 5, each using a different solar analog for calibration. The July 4 data were heavily contaminated by background stars, leaving only one out of six exposures usable. Although the July 4 data have a much lower signal-to-noise ratio (SNR) than those from July 5, the spectral profile and slope are in good agreement across both nights. The measured spectral slopes over the 0.5–0.8 m range are 12.1 1.2 % per 1000 Å (July 4) and 10.1 0.2 % per 1000 Å (July 5). Due to the contamination and limited usable signal, we exclude the July 4 data from further analysis and focus on the higher-quality July 5 dataset.
The red slope in the GMOS data is not continuous; instead, the spectrum shows a slight downturn, resulting in a flattened slope beyond 0.85 m. This downturn may indicate the presence of silicate minerals, which exhibit diagnostic absorption features near 1 m, commonly observed in S-type asteroids. We compared the 3I spectrum with various asteroid taxonomic types and found the best matches to be D-type and L-type asteroids, shown as red and green lines in Figure 1, respectively.
We requested DDT time on the IRTF to follow up on a possible absorption feature beyond 0.9 m, suggested by the shape of the GMOS optical spectrum. The resulting near-infrared spectrum of 3I/ATLAS, obtained with the SpeX instrument, is shown in Figure 2. It maintains the red slope seen in the optical and shows no evidence of silicate-related absorption features near the 1 m region. However, across the 0.9–1.5 m range, the measured slope is 3.06 0.21 % per 1000 Å, noticeably shallower than the optical slope.
Beyond 1.5 m, the spectrum shows a neutral to slightly blue slope of % per 1000 Å over the 1.5–2.3 m range. As indicated in Figure 2, the telluric absorption bands are too narrow and well-calibrated to explain the broad flattening seen between 1.5 and 2.3 m, making it unlikely that atmospheric effects are responsible.
As with the optical data, we compared the near-infrared spectrum to representative asteroid taxonomic types. Consistent with the optical results, the 3I/ATLAS spectrum closely resembles that of D-type asteroids shortward of 1.5 m but diverges from the D-type average at longer wavelengths.
The observed change in spectral slope may be attributed to absorption features from silicates (e.g., pyroxene) or from water ice. However, the absence of a 1 m silicate band and the distinct difference from the S-type spectral profile argue against pyroxene as the cause. We explore the water ice interpretation in the next section.
3.1 Water Ice Detection and Spectral Modeling
We generated the water ice reflectance spectrum using a combination of Mie theory and Hapke radiative transfer modeling, following the approach described in B. Yang et al. (2014). The real part (n) of the refractive index was taken from S. G. Warren (1984), and the imaginary part (k) from W. M. Grundy & B. Schmitt (1998). Based on the heliocentric distance of 3I/ATLAS at the time of observation, we adopted the 120 K dataset to represent appropriate thermal conditions.
As noted in the previous section, the continuum of 3I/ATLAS flattens beyond 1.5 m, a trend that can be explained by the presence of large ice grains. Because the spectral slope of water ice becomes bluer with increasing particle size, we modeled the ice using 10 m grains to account for the near-infrared flattening. This large grain size is consistent with the observed sunward concentration of dust (D. Z. Seligman et al., 2025), as smaller grains would be more strongly affected by solar radiation pressure and shift toward the anti-solar direction.
To model the reddish dust continuum of 3I/ATLAS, we selected material representative of D-type asteroids, as the object’s reflectance below 1.5 m closely resembles that taxonomic class. Previous studies have identified Tagish Lake meteorites as good spectral analogs to D-type asteroids (T. Hiroi et al., 2001). We therefore adopted the reflectance spectrum of a Tagish Lake sample—MT-TXH-024-A, from the RELAB spectral library (C. M. Pieters & T. Hiroi, 2004), to represent the dust component. This material reproduces the moderately red slope observed shortward of 1.5 m and provides a physically motivated baseline for constructing areal mixing models with water ice.
Our best-fit model, shown as the solid red line in Figure 3, consists of an areal mixture of 30% water ice and 70% Tagish Lake material. Despite its simplicity, the model reproduces the 3I/ATLAS spectrum adequately well, capturing both the continuum slope and the broad absorption feature near 2.0 m. We note that the diagnostic 1.5 m absorption band typically associated with water ice is not present in the 3I spectrum. The absence or weakness of the 1.5 μm absorption feature has been observed in the spectra of several active comets containing water ice.(B. Yang et al., 2009, 2014).
The 1.5 m water-ice absorption band is often difficult to detect because it is much shallower than the stronger 2.0 and 3.0 m bands. At low spectral resolution, the 1.5 m feature can be washed out, particularly when the signal-to-noise ratio is low or the spectrum is dominated by coma emission. During the Rosetta mission, the 2.0 and 3.0 m bands were clearly present, while the 1.5 m feature was weak (M. A. Barucci et al., 2016). A similar effect is seen in Centaurs, where water ice is sometimes detected at 2.0 and 3.0 m, but the 1.5 m band remains weak due to mixing with dark surface materials (A. Guilbert et al., 2009).
This simple model provides a reasonable first-order approximation of the coma composition. More sophisticated modeling approaches are beyond the scope of this study and would be better suited to future near-infrared observations of 3I with higher spectral resolution and improved signal quality.
3.2 Comparison to Other Known Objects
We compared the spectrum of 3I/ATLAS with those of other solar system bodies and known ISOs. All three ISOs exhibit moderately red spectral slopes in the optical and the NIR. Among them, 3I has the highest SNR to date, allowing for tighter constraints on its composition and dust properties. While 2I/Borisov shows a much redder slope in the H and K bands, 3I is notably less red in this region, likely due to the presence of water ice, as indicated by the spectral flattening beyond 1.5 m.
The spectrum of 3I also closely resembles that of comet 6P/d’Arrest (B. Yang, 2009), apart from the water ice absorption feature, suggesting similar coma composition. However, both objects are active, and their spectra are dominated by coma dust rather than direct reflections from the nucleus. As such, the observed spectral slope may reflect the particle size distribution in the coma more than the intrinsic surface composition.
Finally, 3I does not exhibit the ultrared optical-NIR colors seen in some trans-Neptunian objects such as Quaoar (D. C. Jewitt & J. X. Luu, 2001), further distinguishing its coma from those of the most compositionally extreme outer solar system bodies.
4 Discussion
As the most abundant solid in protoplanetary disks, water ice plays a key role in planet formation (B. Bitsch & A. Johansen, 2016; J. Müller et al., 2021). In the Solar System, it likely contributed to the cores of giant planets (F. J. Ciesla, 2014) and is a major component of comets (M. F. A’Hearn, 2017; M. J. Mumma & S. B. Charnley, 2011). The detection of water ice in the coma of 3I/ATLAS suggests that it formed beyond the snowline of its parent system, and that water ice could be its major component, similar to comets in the solar system. The current brightening is consistent with activity driven by water-ice sublimation, but since no gas emission has yet been observed, there could also be activity driven by super-volatiles from the sub-surface such as CO or CO2. The H2O/CO ratio remains unknown, making future searches for CO and CO2 with ground- and space-based facilities essential.
The physical structure of water ice records its formation conditions. Below 100 K, water forms thermodynamically unstable amorphous ice, which gradually converts to crystalline ice above 137 K (A. Kouchi et al., 1994). Due to limited SNR in the IRTF data, we cannot constrain the ice phase through spectral modeling. However, the water ice signature appears to be strengthening as the comet’s activity increases. IRTF spectra taken 10 days earlier set a 7% upper limit on coma ice (T. Kareta et al., 2025), while our observations suggest that water ice accounts for approximately 30% of the coma by area. During this period, 3I moved from =4.4 to =4.1 au and brightened significantly. The increased activity not only enhanced the comet’s brightness, improving the signal-to-noise ratio of the spectrum, but may also have promoted the ejection of additional icy grains. Additionally, imperfect telluric correction in the T. Kareta et al. (2025) spectrum may have limited their ability to accurately model the water-ice feature. Upcoming JWST/NIRSpec observations are expected to provide deeper insight into the ice properties and composition.
Beyond the absorption feature near 2.0 m, the spectral slope of the continuum provides key insights into the dust composition of the coma. Several studies have reported optical spectral slopes for 3I/ATLAS using various instruments and observing conditions, with values ranging from 17% to 19% 4% per 1000 Å (C. Opitom et al., 2025; D. Z. Seligman et al., 2025). Similar large ranges of measured spectral slope is also reported among previous studies of 1I and 2I.
Our measurement of = 10% per 1000 Å, derived from Gemini South/GMOS data taken on July 5, is slightly lower than this range. The observed variation in spectral slope could be due to using different solar analog stars. The near-infrared slope measured from IRTF/SpeX over the 0.9–1.5 m range is 3.060.21% per 1000Å, noticeably flatter than the optical slope, consistent with the presence of water ice in the coma.
Despite being obtained 9 days apart with different instruments, the GMOS and SpeX datasets show excellent agreement in the overlapping 0.8–0.9 m region, suggesting temporal stability in the continuum shape and confirming the consistency of the optical–NIR transition. This comparison underscores the importance of coordinated, multi-wavelength observations in characterizing transient interstellar objects. Single-instrument datasets may be affected by instrumental or environmental biases, but combined optical and infrared spectra enable a more robust and comprehensive understanding of composition and activity.
5 Summary
We conducted optical and near-infrared spectroscopic observations of the interstellar object 3I/ATLAS using Gemini South/GMOS on UT 2025 July 5 and IRTF/SpeX on 2025 July 14. Our main findings are as follows:
1.
The near-infrared spectrum shows a neutral to slightly bluish slope of % per 1000 Å over the 1.5-2.3 m range. The flattening beyond 1.5 m and the presence of a broad absorption band centered near 2.0 m are consistent with the presence of water ice grains 10 m in radius, contributing approximately 30% of the coma by area.
2.
The optical spectrum of 3I appears featureless and exhibits a moderately red slope consistent with D-type material of % per 1000 Å, (between 0.5-0.8 m) consistent with the slope range reported in earlier studies.
3.
The GMOS and SpeX datasets, taken with different instruments and separated by 9 days, display consistent spectral slopes in the overlapping 0.8–0.9 m range. This consistency indicates that no significant temporal evolution occurred in the continuum shape during that period.
Together, these results suggest that 3I/ATLAS hosts a coma containing large water ice grains, and that its dust continuum is stable over at least week-long timescales. The spectral characteristics further distinguish 3I from known ultrared trans-Neptunian objects and align it more closely with active Jupiter-family comets.

### v2

@@
 Abstract
-We present optical and near-infrared spectroscopy of the interstellar object 3I/ATLAS, obtained with Gemini-S/GMOS and NASA IRTF/SpeX on 2025 July 5 and July 14.
-The optical spectrum shows a red slope of 10% per 1000 Å between 0.5 and 0.8 m, closely resembling that of typical D-type asteroids.
-At longer wavelengths, the near-infrared spectrum flattens significantly to 3% per 1000 Å from 0.9 to 1.5 m, consistent with the spectral behavior of large water ice grains in the coma.
-Spectral modeling with an areal mixture of 70% Tagish Lake meteorite and 30% 10 m-sized water ice successfully reproduces both the overall continuum and the broad absorption feature near 2.0 m.
-The 1.5 m water ice band, however, is not detected—likely due to the limited signal-to-noise of the IRTF data and dilution by refractory materials.
-The 30% ice fraction should be interpreted as an approximate, order-of-magnitude estimate of the coma composition.
-The agreement between the GMOS and SpeX spectra, taken nine days apart, suggests short-term stability in the coma’s optical properties.
-Our observations reveal that 3I/ATLAS is an active interstellar comet containing abundant water ice, with a dust composition more similar to D-type asteroids than to ultrared trans-Neptunian objects.
-UTF8gbsn
+We present optical and near-infrared spectroscopy of the interstellar object 3I/ATLAS, obtained with Gemini-S/GMOS and NASA IRTF/SpeX on 2025 July 5 and 14.
+The optical spectrum shows a red slope of 11% per 1000 Å between 0.5 and 0.8 m, resembling typical D-type asteroids and distinct from ultrared trans-Neptunian objects.
+At longer wavelengths, the near-infrared continuum flattens to 3% per 1000 Å between 0.9 and 1.5 m, with a broad absorption feature near 2.0 m indicative of water ice grains in the coma.
+Spectral modeling with a mixture of 63% amorphous carbon and 37% 1 m-sized water ice reproduces both the continuum and the 2.0 m band, while the 1.5 m water ice band is not detected, likely due to limited signal-to-noise in the IRTF data and dilution by refractory material.
+The close agreement between the GMOS and SpeX spectra, taken nine days apart, indicates short-term stability in the coma’s optical properties.
+These observations demonstrate that 3I/ATLAS is an active interstellar comet containing abundant water ice, consistent with the theoretical expectation that its home planetary system had a high bulk fraction of water ice by mass.
 1 Introduction
@@
 Tonry et al., 2018) during its robotic sky survey and was initially designated A11pl3Z.
-This was a particularly challenging detection because the object lay in Sagittarius, near the galactic plane, a region often avoided by most surveys due to stellar crowding.
-Following the discovery, numerous observatories quickly conducted follow-up observations, and archival searches uncovered pre-discovery images that extended the orbital arc back to 2025 May 22.
-Early measurements indicated an orbital eccentricity greater than 6, and reports of coma or activity prompted its official designation as 3I/ATLAS = C/2025 N1 (ATLAS) (MPEC 2025 N12).
-The initial characterization of 3I/ATLAS (D.
+This was a particularly challenging detection because the object lay in Sagittarius, near the galactic plane, a region often avoided by most surveys due to stellar crowding (Z.
+T.
+F.
+Collaboration et al., 2025).
+Following the discovery, numerous observatories quickly conducted follow-up observations, and archival searches uncovered pre-discovery images that extended the orbital arc back to 2025 May 22 (A.
+D.
+Feinstein et al., 2025).
+Early measurements indicated an orbital eccentricity greater than 6, and reports of coma or activity prompted its official designation as 3I/ATLAS = C/2025 N1 (ATLAS).
+The initial characterization studies of 3I/ATLAS (D.
 Z.
-Seligman et al., 2025) highlighted both similarities and differences relative to the first two known interstellar objects.
+Seligman et al., 2025; R.
+de la Fuente Marcos et al., 2025; B.
+T.
+Bolin et al., 2025), highlighted both similarities and differences relative to the first two known interstellar objects.
 Whereas 1I/ʻOumuamua and 2I/Borisov were both estimated to have sub-kilometer radii (A.
-Fitzsimmons et al., 2024), early photometric measurements for 3I suggested a nucleus radius of less than 10 km.
-This estimate was derived from scattered light measurements that included contributions from both the nucleus and surrounding coma dust, assuming a typical cometary albedo.
+Fitzsimmons et al., 2024), early imaging with the Hubble Space Telescope constrains the nucleus diameter to 0.44–5.6 km (D.
+Jewitt et al., 2025).
 Unlike 1I/ʻOumuamua, which showed no detectable dust coma during its brief observation window, both 2I/Borisov and 3I/ATLAS were discovered with prominent dust comae.
@@
 de la Fuente Marcos et al., 2025).
-Spectral reflectance measurements of 3I consistently indicate red dust, with reported spectral slopes of = 17–19% per 1000 Å between 0.4–0.8 m (D.
+James Webb Space Telescope (JWST) Near Infrared Spectrograph (NIRSpec) spectroscopy at shows a -dominated coma with , and detections of , CO, OCS, water ice, and dust (M.
+A.
+Cordiner et al., 2025).
+Swift/UVOT detected OH near , giving at , consistent with activity from icy grains (Z.
+Xing et al., 2025).
+Optical spectra obtained two days after discovery showed a red, nearly featureless reflectance, with no strong gas bands at that time, indicate red dust, with reported spectral slopes of = 17–19% per 1000 Å between 0.4–0.8 m(C.
+Opitom et al., 2025; D.
 Z.
-Seligman et al., 2025; C.
-Opitom et al., 2025; M.
-Belyakov et al., 2025), similar to D-type asteroids, and a flattening of the slope beyond 0.8 m.
-Gas emissions from CN were detected for 2I/Borisov about a week after discovery (A.
-Fitzsimmons et al., 2019) at production rate of Q(CN) = 4 1024 molec/s when it was at =2.7 au.
-In contrast, to date, no gas emissions have been detected from 3I with an upper limit of Q(CN)4.91023 molec/s (A.
-Alvarez-Candal et al., 2025).
-However, this limit was obtained at 4.4 au where it is still quite cold to expect CN.
-Only a handful of comets have ever had CN detected beyond 4 au and on the order of a dozen beyond 3 au (M.
-F.
-A’Hearn et al., 1995; A.
-L.
-Cochran et al., 2012).
+Seligman et al., 2025) later VLT/UVES and X-shooter data revealed emergent CN emission and numerous Ni I lines at (R.
+Rahatgaonkar et al., 2025).
+An assessment of 3I/ATLAS within the Ōtautahi–Oxford model infers, from its kinematics, an age exceeding 7.6 Gyr and a high H2O mass fraction (M.
+J.
+Hopkins et al., 2025).
 In this paper, we present spectroscopic observations of 3I/ATLAS aimed at characterizing its dust coma.
@@
 2.1 Gemini
-The Gemini Multi-Object Spectrograph (GMOS) on the Gemini South telescope was used to obtain low resolution spectra of 3I/ATLAS on 2025 July 4 and 5 UT.
+The Gemini Multi-Object Spectrograph (GMOS) on the Gemini South telescope (D.
+Crampton et al., 2000) was used to obtain low resolution spectra of 3I/ATLAS on 2025 July 4 and 5 UT.
 Triggering a rapid target of opportunity program [GS-2025A-Q-113], we obtained 6 exposures using the R150 grating and a 1.0′′ slit, with central wavelength at 0.6 m.
 We used a spatial binning of 22 and spectral binning 44.
-We also also observed two G2V solar analog stars (HD 174466 & HD 167221) at airmasses similar to 3I, serving as both telluric correction standards and as solar analog references.
+We also observed two G2V solar analog stars (HD 174466 & HD 167221) at airmasses similar to 3I, serving as both telluric correction standards and as solar analog references.
+GMOS data were calibrated with bias frames, internal spectroscopic flats, and CuAr arc lamp exposures taken close in time to the science frames for the wavelength solution.
 The data were reduced using the DRAGONS (Data Reduction for Astronomy from Gemini Observatory North and South) pipeline (K.
 Labrie et al., 2019), the official data reduction package provided by the Gemini Observatory.
-| Observation Date | Tel./Inst. | Object | Exposure | Airmass | aaHeliocentric distance | bbGeocentric distance | ccPhase angle |
-| (UT) | (s) | (au) | (au) | (∘) | |||
+| Observation Date | Tel./Inst. | Object$\dagger$$\dagger$footnotemark: | Exposure | Airmass | aaHeliocentric distance | bbGeocentric distance | (∘) |
+| (UT) | (s) | (au) | (au) | Phase | |||
 | 2025-07-04 | Gemini/GMOS | 3I | 6200 | 1.26–1.35 | 4.404 | 3.408 | 2.96 |
@@
 2.2 IRTF
-Observations were conducted on the evening of 2025 July 14 (UT) using Director’s Discretionary Time (DDT) with the SpeX spectrograph on the NASA Infrared Telescope Facility (IRTF).
+Observations were conducted on the evening of 2025 July 14 (UT) using Director’s Discretionary Time (DDT) with the SpeX spectrograph (J.
+T.
+Rayner et al., 2003) on the NASA Infrared Telescope Facility (IRTF).
 We conducted observations using the SpeX instrument in combination with the MIT Optical Rapid Imaging System (MORIS) camera (A.
@@
 This setup enabled stable guiding of the faint interstellar object in a crowded field, while simultaneously acquiring near-infrared spectra with SpeX.
-NIR data were acquired in prism mode covering the 0.8–2.5 m wavelength range, using a 0.8′′ 15′′ slit, which provided a spectral resolving power of R150 (J.
+NIR data were acquired in prism mode covering the 0.8–2.5 m wavelength range, using a 0.8′′ 15′′ slit, which provided a spectral resolving power of R100 (J.
 T.
 Rayner et al., 2003).
-Two nearby G-type solar analog star observed along with 3I, served both as the telluric correction standard and as a solar analog reference.
+Two nearby G2V solar analog star observed along with 3I, served both as the telluric correction standard and as a solar analog reference.
+Both the science target and the solar analogs were observed at the parallactic angle.
+Following each IRTF/SpeX science sequence, we obtained internal flat fields and arc lamp exposures for wavelength calibration.
 The spectral data were reduced using the SpeXtool data reduction pipeline (M.
@@
 3 Observational Analysis
-The GMOS spectrum of 3I/ATLAS is shown in Figure 1.
+The GMOS spectrum of 3I/ATLAS, normalized at 0.55 m, is shown in Figure 1.
 It is largely featureless and exhibits a moderately red spectral slope.
@@
 Although the July 4 data have a much lower signal-to-noise ratio (SNR) than those from July 5, the spectral profile and slope are in good agreement across both nights.
-The measured spectral slopes over the 0.5–0.8 m range are 12.1 1.2 % per 1000 Å (July 4) and 10.1 0.2 % per 1000 Å (July 5).
+The measured spectral slopes over the 0.5–0.8 m range are 14.1 1.3 % per 1000 Å (July 4) and 11.4 0.2 % per 1000 Å (July 5).
 Due to the contamination and limited usable signal, we exclude the July 4 data from further analysis and focus on the higher-quality July 5 dataset.
-The red slope in the GMOS data is not continuous; instead, the spectrum shows a slight downturn, resulting in a flattened slope beyond 0.85 m.
-This downturn may indicate the presence of silicate minerals, which exhibit diagnostic absorption features near 1 m, commonly observed in S-type asteroids.
-We compared the 3I spectrum with various asteroid taxonomic types and found the best matches to be D-type and L-type asteroids, shown as red and green lines in Figure 1, respectively.
+To place 3I/ATLAS in the context of known Solar System populations, we compared its reflectance spectrum to representative asteroid taxonomic classes and to ultrared Kuiper belt objects (D.
+C.
+Jewitt, 2002), shown in Figure 1.
+We note that because the spectrum of 3I is coma dominated, the spectral slope need not reflect composition alone; it is also shaped by the grain size distribution in the coma.
+We found that the 3I spectrum is broadly similar to that of D-type asteroids with a moderately red spectral slope, but it does not exhibit the ultrared optical-NIR colors seen in some trans-Neptunian objects such as Varuna, further distinguishing its coma from those of the most compositionally extreme outer solar system bodies.
+However, the red slope in the GMOS data is not continuous; instead, the spectrum shows a slight downturn, resulting in a flattened slope beyond 0.85 m.
+This downturn may indicate the presence of silicate minerals, which exhibit diagnostic absorption features near 1m, commonly observed in S-type asteroids, shown as the green line in Figure 1.
 We requested DDT time on the IRTF to follow up on a possible absorption feature beyond 0.9 m, suggested by the shape of the GMOS optical spectrum.
+Similar to the previous GMOS observations, two nearby G2V stars were observed as solar analogs.
+The resulting reflectance spectra show comparable profiles.
+However, the airmass difference between 3I and HD 156734 was larger, leading to a poorer telluric correction.
+We therefore used HD 167221 to compute the reflectance and for subsequent analysis.
 The resulting near-infrared spectrum of 3I/ATLAS, obtained with the SpeX instrument, is shown in Figure 2.
-It maintains the red slope seen in the optical and shows no evidence of silicate-related absorption features near the 1 m region.
-However, across the 0.9–1.5 m range, the measured slope is 3.06 0.21 % per 1000 Å, noticeably shallower than the optical slope.
-Beyond 1.5 m, the spectrum shows a neutral to slightly blue slope of % per 1000 Å over the 1.5–2.3 m range.
-As indicated in Figure 2, the telluric absorption bands are too narrow and well-calibrated to explain the broad flattening seen between 1.5 and 2.3 m, making it unlikely that atmospheric effects are responsible.
-As with the optical data, we compared the near-infrared spectrum to representative asteroid taxonomic types.
-Consistent with the optical results, the 3I/ATLAS spectrum closely resembles that of D-type asteroids shortward of 1.5 m but diverges from the D-type average at longer wavelengths.
-The observed change in spectral slope may be attributed to absorption features from silicates (e.g., pyroxene) or from water ice.
-However, the absence of a 1 m silicate band and the distinct difference from the S-type spectral profile argue against pyroxene as the cause.
-We explore the water ice interpretation in the next section.
+To compare the GMOS data with the IRTF spectrum, we re-normalized the GMOS to 0.85 m and normalized the IRTF spectrum at the same wavelength.
+In the overlapping region across the 0.7–0.9 m range, the measured GMOS slope is 5.2 0.2 % per 1000 Å and the IRTF slope is 6.7 1.4 % per 1000 Å.
+Beyond 0.9 m, the GMOS spectrum shows a negative spectral slope of % per 1000,Å, whereas the IRTF spectrum maintains the positive slope observed in the optical.
+To investigate this discrepancy, we compared our data with X-shooter/VLT observations (A.
+Alvarez-Candal et al., 2025) obtained just one day prior to the GMOS dataset.
+The X-shooter spectrum confirms the continuation of a positive slope from 0.9 to 1.2 m, with no evidence of silicate-related absorption features near 1.0 m.
+The apparent negative slope arises at the red edge of the GMOS spectrum, where the instrument sensitivity is lowest; consequently, this portion of the data is unreliable and cannot be interpreted as intrinsic to the object.
+To better illustrate the continuum trend, we overplot a 40-pixel smoothed spectrum in Figure 2.
+The spectrum of 3I exhibits a broad but shallow absorption feature centered at 2.05 m (marked by a red arrow) with a depth of %.
+Such a feature could in principle be attributed to silicates, which commonly produce absorptions near 2.0 m in S-type asteroids (shown in blue).
+However, the absence of the characteristic 1.0 m silicate band and the clear mismatch with the S-type spectral profile make a pyroxene origin unlikely.
+An alternative explanation is water ice, the dominant volatile in most Solar System comets; we examine this possibility in the following section.
+In addition, a narrow absorption feature is observed between 2.2 and 2.3 m (marked by a black arrow), though its origin cannot be unambiguously identified.
+One possibility is that it may arise from C–H vibrations in organic or carbonaceous materials.
+Additional observations will be necessary to verify the presence of this feature and better constrain its source.
 3.1 Water Ice Detection and Spectral Modeling
-We generated the water ice reflectance spectrum using a combination of Mie theory and Hapke radiative transfer modeling, following the approach described in B.
-Yang et al.
-(2014).
-The real part (n) of the refractive index was taken from S.
+We generated a synthetic spectral model with amorphous carbon and water ice, computed using Mie theory and following the approach described in R. Zhao et al. (2025).
+The real part () of the refractive index of water ice was taken from S.
 G.
-Warren (1984), and the imaginary part (k) from W.
+Warren (1984), and the imaginary part () from W.
 M.
@@
 Schmitt (1998).
-Based on the heliocentric distance of 3I/ATLAS at the time of observation, we adopted the 120 K dataset to represent appropriate thermal conditions.
-As noted in the previous section, the continuum of 3I/ATLAS flattens beyond 1.5 m, a trend that can be explained by the presence of large ice grains.
-Because the spectral slope of water ice becomes bluer with increasing particle size, we modeled the ice using 10 m grains to account for the near-infrared flattening.
-This large grain size is consistent with the observed sunward concentration of dust (D.
-Z.
-Seligman et al., 2025), as smaller grains would be more strongly affected by solar radiation pressure and shift toward the anti-solar direction.
-To model the reddish dust continuum of 3I/ATLAS, we selected material representative of D-type asteroids, as the object’s reflectance below 1.5 m closely resembles that taxonomic class.
-Previous studies have identified Tagish Lake meteorites as good spectral analogs to D-type asteroids (T.
-Hiroi et al., 2001).
-We therefore adopted the reflectance spectrum of a Tagish Lake sample—MT-TXH-024-A, from the RELAB spectral library (C.
+This approach of combining two different sets of optical constants has been adopted in earlier studies (e.g., K.
+D.
+Seelos et al., 2008; W.
 M.
-Pieters & T.
-Hiroi, 2004), to represent the dust component.
-This material reproduces the moderately red slope observed shortward of 1.5 m and provides a physically motivated baseline for constructing areal mixing models with water ice.
-Our best-fit model, shown as the solid red line in Figure 3, consists of an areal mixture of 30% water ice and 70% Tagish Lake material.
-Despite its simplicity, the model reproduces the 3I/ATLAS spectrum adequately well, capturing both the continuum slope and the broad absorption feature near 2.0 m.
-We note that the diagnostic 1.5 m absorption band typically associated with water ice is not present in the 3I spectrum.
-The absence or weakness of the 1.5 μm absorption feature has been observed in the spectra of several active comets containing water ice.(B.
-Yang et al., 2009, 2014).
-The 1.5 m water-ice absorption band is often difficult to detect because it is much shallower than the stronger 2.0 and 3.0 m bands.
-At low spectral resolution, the 1.5 m feature can be washed out, particularly when the signal-to-noise ratio is low or the spectrum is dominated by coma emission.
-During the Rosetta mission, the 2.0 and 3.0 m bands were clearly present, while the 1.5 m feature was weak (M.
+Grundy, 2009).
+Based on the heliocentric distance of 3I/ATLAS at the time of observation, and assuming an albedo of 0.5 for pure water ice and an emissivity of 0.9, the equilibrium temperature of ice grains at 4 au is about 120 K.
+We verified the robustness of our results by testing alternative optical constant datasets, and found the differences in the modeled spectra to be negligible.
+We did not choose the widely applied dataset of R.
+M.
+Mastrapa et al.
+(2009), as those optical constants begin at 1.1 m and therefore omit the shorter wavelength range still valuable for our analysis.
+The optical constants of amorphous carbon, used as a representative refractory material, were taken from (F.
+Rouleau & P.
+G.
+Martin, 1991).
+We assumed that each of the two compositions follows a modified power-law differential size distribution over the range of , i.e.,
+where is the power-law index, and is related to the peak size by .
+Our best-fit model, shown as the solid red line in Figure 3, successfully reproduces the 3I/ATLAS spectrum well, capturing both the continuum slope and the broad absorption feature near 2.0 m.
+It consists of a volume mixture of 63% amorphous carbon and 37% water ice at a temperature of 120 K.
+The carbon grain size distribution follows a power-law with an index of 3.6, peaking at 0.10 m, which is much smaller than the peak size of the water ice grains at 1.3 m.
+The water ice grain size distribution is characterized by a steeper power-law index of 4.6, indicating a relatively larger population of small grains.
+We note that, given the limited signal-to-noise ratio (SNR) of the IRTF spectrum and the broad parameter space, our best-fit solution is not unique; other combinations of water-ice fraction and grain-size distribution can yield acceptable, though not optimal, fits to the data.
+But in any case, the presence of water ice is required to reproduce the absorption feature near 2.0 m.
+The diagnostic 1.5 m absorption band typically associated with water ice is absent from the 3I spectrum.
+Such suppression of the 1.5 m band has also been reported in the spectra of several active comets known to contain water ice (B.
+Yang et al., 2009, 2014, also see Figure 4).
+During the Rosetta mission, the 2.0 and 3.0 m bands were clearly present, while the 1.5 m band was weak (M.
 A.
@@
 Guilbert et al., 2009).
-This simple model provides a reasonable first-order approximation of the coma composition.
-More sophisticated modeling approaches are beyond the scope of this study and would be better suited to future near-infrared observations of 3I with higher spectral resolution and improved signal quality.
+The 1.5 m water-ice absorption band is often difficult to detect because it is much shallower than the stronger 2.0 and 3.0 m bands.
+At low spectral resolution, the 1.5 m feature can be washed out, particularly when the SNR is low or the spectrum is dominated by coma emission.
+As shown by B.
+Yang et al.
+(2014), particle size strongly influences the strength of water-ice absorption features.
+In our spectral model, relatively small water-ice grains produce only a weak 1.5 m band, which could easily be obscured by the noise in the IRTF data.
 3.2 Comparison to Other Known Objects
@@
 While 2I/Borisov shows a much redder slope in the H and K bands, 3I is notably less red in this region, likely due to the presence of water ice, as indicated by the spectral flattening beyond 1.5 m.
-The spectrum of 3I also closely resembles that of comet 6P/d’Arrest (B.
-Yang, 2009), apart from the water ice absorption feature, suggesting similar coma composition.
-However, both objects are active, and their spectra are dominated by coma dust rather than direct reflections from the nucleus.
-As such, the observed spectral slope may reflect the particle size distribution in the coma more than the intrinsic surface composition.
-Finally, 3I does not exhibit the ultrared optical-NIR colors seen in some trans-Neptunian objects such as Quaoar (D.
-C.
-Jewitt & J.
-X.
-Luu, 2001), further distinguishing its coma from those of the most compositionally extreme outer solar system bodies.
+The spectrum of 3I closely resembles that of the Oort Cloud comet C/2006 W3 (B.
+Yang, 2009), which likewise lacks the 1.5 m absorption feature but shows a strong 2.0 m water-ice band.
+In addition, the optical–NIR spectral slope of 3I is very similar to that of the Jupiter-family comet 6P/d’Arrest (B.
+Yang, 2009).
+Despite their distinct source regions, 3I and these active Solar System comets exhibit striking spectral similarities.
+Because the scattering cross section of the dust coma dominates over that of the small nucleus, the observed resemblance likely reflects comparable particle-size distributions or similar dust compositions in their comae.
 4 Discussion
@@
 The detection of water ice in the coma of 3I/ATLAS suggests that it formed beyond the snowline of its parent system, and that water ice could be its major component, similar to comets in the solar system.
-The current brightening is consistent with activity driven by water-ice sublimation, but since no gas emission has yet been observed, there could also be activity driven by super-volatiles from the sub-surface such as CO or CO2.
-The H2O/CO ratio remains unknown, making future searches for CO and CO2 with ground- and space-based facilities essential.
+The derived abundance of 37% water ice by volume in the coma is consistent with the theoretical study predicting a water-ice mass fraction of 20–40% for 3I-like objects (M.
+J.
+Hopkins et al., 2025).
 The physical structure of water ice records its formation conditions.
@@
 Due to limited SNR in the IRTF data, we cannot constrain the ice phase through spectral modeling.
-However, the water ice signature appears to be strengthening as the comet’s activity increases.
 IRTF spectra taken 10 days earlier set a 7% upper limit on coma ice (T.
-Kareta et al., 2025), while our observations suggest that water ice accounts for approximately 30% of the coma by area.
-During this period, 3I moved from =4.4 to =4.1 au and brightened significantly.
-The increased activity not only enhanced the comet’s brightness, improving the signal-to-noise ratio of the spectrum, but may also have promoted the ejection of additional icy grains.
-Additionally, imperfect telluric correction in the T.
-Kareta et al.
-(2025) spectrum may have limited their ability to accurately model the water-ice feature.
-Upcoming JWST/NIRSpec observations are expected to provide deeper insight into the ice properties and composition.
+Kareta et al., 2025), while our observations suggest that water ice accounts for approximately 37% of the coma by volume.
+In addition, recent IRTF observations show that the water ice signature appears to be stronger as the comet’s activity increases (C.
+M.
+Lisse et al., 2025).
+The increased activity not only enhanced the comet’s brightness, improving the SNR of the spectrum, but may also have promoted the ejection of additional icy grains.
+The JWST/NIRSpec observations clearly show strong water-ice absorption features at 3.0 and 4.5 m (M.
+A.
+Cordiner et al., 2025).
+Detailed spectral modeling of these data can provide valuable information on the size, temperature, and impurities of the water-ice grains, and place strong constraints on the formation conditions of this ISO.
 Beyond the absorption feature near 2.0 m, the spectral slope of the continuum provides key insights into the dust composition of the coma.
-Several studies have reported optical spectral slopes for 3I/ATLAS using various instruments and observing conditions, with values ranging from 17% to 19% 4% per 1000 Å (C.
+Several studies have reported optical spectral slopes for 3I/ATLAS using various instruments and observing conditions, with values ranging from 16% to 27% per 1000 Å (C.
 Opitom et al., 2025; D.
 Z.
-Seligman et al., 2025).
+Seligman et al., 2025; T.
+H.
+Puzia et al., 2025; B.
+T.
+Bolin et al., 2025).
 Similar large ranges of measured spectral slope is also reported among previous studies of 1I and 2I.
-Our measurement of = 10% per 1000 Å, derived from Gemini South/GMOS data taken on July 5, is slightly lower than this range.
+Our measurement of = 11% per 1000 Å, derived from Gemini South/GMOS data taken on July 5, is lower than this range.
 The observed variation in spectral slope could be due to using different solar analog stars.
-The near-infrared slope measured from IRTF/SpeX over the 0.9–1.5 m range is 3.060.21% per 1000Å, noticeably flatter than the optical slope, consistent with the presence of water ice in the coma.
-Despite being obtained 9 days apart with different instruments, the GMOS and SpeX datasets show excellent agreement in the overlapping 0.8–0.9 m region, suggesting temporal stability in the continuum shape and confirming the consistency of the optical–NIR transition.
+The near-infrared slope measured with IRTF/SpeX over the 0.8–2.3 m range is % per 1000 Å, significantly flatter than the optical slope, and consistent with the presence of water ice in the coma.
+Despite being obtained 9 days apart with different instruments, the GMOS and SpeX datasets are statistically consistent in the overlapping 0.7–0.9 m region, indicating temporal stability in the continuum shape and confirming the reliability of the optical-NIR transition.
 This comparison underscores the importance of coordinated, multi-wavelength observations in characterizing transient interstellar objects.
@@
 1.
-The near-infrared spectrum shows a neutral to slightly bluish slope of % per 1000 Å over the 1.5-2.3 m range.
-The flattening beyond 1.5 m and the presence of a broad absorption band centered near 2.0 m are consistent with the presence of water ice grains 10 m in radius, contributing approximately 30% of the coma by area.
+The optical spectrum of 3I/ATLAS, obtained with GMOS, is featureless and exhibits a moderately red slope of % per 1000 Å between 0.5 and 0.8 m.
+This spectral behavior is comparable to D-type asteroids and active comets in the Solar System, but markedly different from the ultrared objects in the trans-Neptunian region.
+The measurement is consistent with slopes reported in earlier studies and supports the interpretation that 3I’s dust is similar in reflectance properties to Solar System cometary material rather than the ultrared TNO population.
 2.
-The optical spectrum of 3I appears featureless and exhibits a moderately red slope consistent with D-type material of % per 1000 Å, (between 0.5-0.8 m) consistent with the slope range reported in earlier studies.
+The near-infrared spectrum from IRTF/SpeX displays a neutral to slightly bluish slope between 1.7 and 2.3 m, together with a broad 11% absorption band centered near 2.0 m.
+These features are well reproduced by a spectral model consisting of amorphous carbon mixed with 1.3 m-sized water ice grains, with water ice contributing about 37% of the coma volume.
+The flatter continuum relative to the optical, combined with the detection of the 2.0 m absorption, provides strong evidence that abundant water ice is present in the coma.
 3.
-The GMOS and SpeX datasets, taken with different instruments and separated by 9 days, display consistent spectral slopes in the overlapping 0.8–0.9 m range.
-This consistency indicates that no significant temporal evolution occurred in the continuum shape during that period.
-Together, these results suggest that 3I/ATLAS hosts a coma containing large water ice grains, and that its dust continuum is stable over at least week-long timescales.
-The spectral characteristics further distinguish 3I from known ultrared trans-Neptunian objects and align it more closely with active Jupiter-family comets.
+The GMOS and SpeX datasets, taken with different instruments and separated by nine days, display consistent spectral slopes in the overlapping 0.7–0.9 m region.
+This agreement indicates that no significant short-term spectral evolution occurred during that period, the overall scattering properties of the coma remained stable over timescales of at least one week.
+Our detection of H2O ice in 3I/ATLAS at 4 au with IRTF, together with the strong CO2 outgassing revealed by JWST and SPHEREx, supports the hypothesis that this object formed in a cold, volatile-rich region of its parent protoplanetary disk.
+These observations also demonstrate that such volatile compositions can be preserved through ejection and maintained during the subsequent interstellar journey.